Skip to main contentSkip to solution

Let XX denote the number of hours you play during a randomly selected day. The probability that XX can take values xx has the following form, where cc is some constant.

P(X=x)={0.1, if x=0cx, if x=1 or x=2c(5−x), if x=3 or x=40, otherwise \mathrm{P}(\mathrm{X}=\mathrm{x})=\left\{\begin{array}{lll} 0.1, & \text { if } \mathrm{x}=0 \\ \mathrm{cx}, & \text { if } \mathrm{x}=1 \text { or } \mathrm{x}=2 \\ \mathrm{c}(5-\mathrm{x}), & \text { if } \mathrm{x}=3 \text { or } \mathrm{x}=4 \\ 0, & \text { otherwise } \end{array}\right.

Match List-I with List-II :

List-IList-II
(A) cc(I) 0.75
(B) P(X≤2)P(X \leq 2)(II) 0.3
(C) P(X=2)P(X = 2)(III) 0.55
(D) P(X≥2)P(X \geq 2)(IV) 0.15

Solution

✅ Correct Option: 2

Let's find the value of constant cc.

Since this is a probability distribution, the sum of all probabilities must equal 1.

P(X=0)+P(X=1)+P(X=2)+P(X=3)+P(X=4)=1P(X = 0) + P(X = 1) + P(X = 2) + P(X = 3) + P(X = 4) = 1

0.1+c(1)+c(2)+c(5−3)+c(5−4)=10.1 + c(1) + c(2) + c(5-3) + c(5-4) = 1

0.1+c+2c+2c+c=10.1 + c + 2c + 2c + c = 1

0.1+6c=10.1 + 6c = 1

6c=0.96c = 0.9

c=0.15c = 0.15


Now let's calculate each probability:

P(X=0)=0.1P(X = 0) = 0.1

P(X=1)=c(1)=0.15P(X = 1) = c(1) = 0.15

P(X=2)=c(2)=0.3P(X = 2) = c(2) = 0.3

P(X=3)=c(5−3)=0.3P(X = 3) = c(5-3) = 0.3

P(X=4)=c(5−4)=0.15P(X = 4) = c(5-4) = 0.15


Let's find the requested probabilities:

(A) c=0.15c = 0.15

(B) P(X≤2)=P(X=0)+P(X=1)+P(X=2)=0.1+0.15+0.3=0.55P(X \leq 2) = P(X = 0) + P(X = 1) + P(X = 2) = 0.1 + 0.15 + 0.3 = 0.55

(C) P(X=2)=0.3P(X = 2) = 0.3

(D) P(X≥2)=P(X=2)+P(X=3)+P(X=4)=0.3+0.3+0.15=0.75P(X \geq 2) = P(X = 2) + P(X = 3) + P(X = 4) = 0.3 + 0.3 + 0.15 = 0.75


Therefore, the matching is:

(A) cc → (IV) 0.15

(B) P(X≤2)P(X \leq 2) → (III) 0.55

(C) P(X=2)P(X = 2) → (II) 0.3

(D) P(X≥2)P(X \geq 2) → (I) 0.75

Keyboard Shortcuts

  • Left arrow: Previous question
  • Right arrow: Next question
  • S key: Jump to solution
  • Q key: Jump to question